# Ytterbium atom interferometry for dark matter searches

### v1

@@ Abstract
Fig. 6 shows the projected sensitivity of atom-interferometric acceleration measurements to vector dark matter as a function of the dark matter particle mass. For the JHU projections, we assume the same experimental parameters as in Section III.A ($n=500$, maximum wave packet separation $2$ m). Likewise, the MAGIS-100 projections assume $n=500$ and $T=4.5$ s. Existing constraints from the MICROSCOPE Touboul et al. (2022), Eöt-Wash Shaw et al. (2022), and LIGO/VIRGO experiments Abbott (2022) are also shown. The searches for oscillating dark matter take into account the stochastic amplitude of the dark matter field Centers et al. (2021). In addition, these searches lose sensitivity at high frequencies when the period of the dark matter oscillation becomes shorter than the coherence time of a single measurement. Sensitivity at high frequencies can be enhanced by resonant detection schemes Graham et al. (2016b), which we do not consider here.
@@
The strongest searches for vector dark matter are expected to be derived from static equivalence principle tests. In this operating mode, the JHU apparatus is expected to reach a similar sensitivity to the MICROSCOPE experiment, while an analogous test in MAGIS-100 would be more sensitive by a factor of 10. Although the sensitivity of searches for time-oscillating forces is generally lower, we note that a search for oscillatory dark matter with Yb accelerometry in MAGIS-100 would reach a comparable sensitivity to MICROSCOPE and would be technically simpler than a static equivalence principle test.
III.3 C. Axion dark matter
@@
Pseudoscalar particles (axions) are well-motivated dark matter candidates because they appear in theories that attempt to resolve other outstanding issues in fundamental physics, such as the strong CP problem Peccei and Quinn (1977) and the hierarchy problem Graham et al. (2015). Models inspired by string theory Arvanitaki et al. (2010) also predict the existence of low-mass axions. Here we consider the detection of axions by means of the spin torque that they exert on nucleons. In the presence of an axion field, a nuclear spin $\mathbf{s}$ experiences a Hamiltonian Graham et al. (2018)
@@
where $g_{\text{aNN}}$ is the coupling constant, the axion field $a$ is given by
@@
the magnitude of the field $a_{0}=\sqrt{2\rho_{\text{DM}}}/(m_{\text{DM}}c)$, and the axion momentum is $\mathbf{p}$. This interaction, which has the same form as a magnetic interaction, gives rise to an energy shift between states with opposite spin direction. Assuming that the axion momentum is determined by its virial velocity, the energy shift is given by Graham et al. (2018)
@@
Figure 7: Interferometer geometry for searching for spin torque due to axion dark matter. Doppler-free two-photon transitions at time $t=0$ create superpositions of the ${}^{1}S_{0}$ electronic ground state and the ${}^{3}P_{0}$ excited state. The differential phase between the $\Delta m=1$ interferometer (a) and the $\Delta m=0$ interferometer (b) is sensitive to spin torque but insensitive to the electronic transition frequency. To reduce systematic effects, the two interferometers are spatially overlapped; interferometers are spaced vertically in the diagram for clarity.
@@
Figure 8: Alternative interferometer geometry for searching for spin torque due to axion dark matter. In (a), the atoms are initialized in the electronic ground state, and a Ramsey interferometer is implemented between hyperfine ground state levels. In (b), the Ramsey interferometer is instead carried out between magnetically sensitive states $\left|M\uparrow\right>$ and $\left|M\downarrow\right>$ (e.g. in the metastable ${}^{3}P_{2}$ level). The interferometer in (a) is sensitive to spin torque but relatively insensitive to the magnetic field, while the interferometer in (b) measures the magnetic field to characterize its systematic effect.
@@
Atom interferometers can be sensitive to this axion dark matter coupling by searching for time dependence in the frequency of any transition that flips a nucleon spin. Following the proposal of Ref. Graham et al. (2018), an interferometer geometry that could be used for such a measurement is shown in Fig. 7. This configuration is sensitive to spin torque via the interferometer with the $\Delta m=1$ transition, while the interferometer with the $\Delta m=0$ transition is used to suppress systematic effects such as laser frequency drift. An alternative approach is illustrated in Fig. 8. Here the Ramsey interferometer between the hyperfine ground states of 171Yb is sensitive to the axion field, while the second Ramsey interferometer between magnetically sensitive states (e.g. states in the metastable ${}^{3}P_{2}$ manifold) is used as a comagnetometer to decorrelate phase shifts from the magnetic field. We note that 171Yb, which has nuclear angular momentum $I=1/2$ from the spin of a single unpaired neutron, provides a simpler platform for this measurement than 87Sr Graham et al. (2018), which has $I=9/2$ from a combination of spin and orbital angular momentum.
@@
Figure 9: Constraint plot and projected sensitivities to axion dark matter coupled to nuclear spin. JHU (ST): projected sensitivity of spin torque search in JHU apparatus with interferometer geometry shown in Fig. 7 or Fig. 8. MAGIS-100 (ST): projected sensitivity of spin torque search with Yb in MAGIS-100. Also shown are existing constraints from atomic magnetometry experiments (Bloch et al. Bloch et al. (2020), green curve; Lee et al. Lee et al. (2023), purple curve). Shaded regions are excluded by existing constraints. All curves are 95% confidence limits.
@@
The projected sensitivity of the JHU apparatus to axion dark matter through the axion-nucleon coupling is shown in Fig. 9. We also plot existing bounds from atomic magnetometer experiments Bloch et al. (2020); Lee et al. (2023) and the projected sensitivity of a spin torque search utilizing 171Yb in MAGIS-100, assuming the same phase resolution as in the JHU apparatus and an interferometer time of $T=4.5$ s. Although the laboratory-scale interferometer is not expected to be competitive with the magnetometer limits, a long-baseline interferometric measurement would have comparable sensitivity in several regions of parameter space.
@@
The spectroscopic search for scalar dark matter [Figs. 1 and 2; sensitivity curve “JHU $\alpha(t)$” in Fig. 4] relies on a differential frequency measurement between the ${}^{1}S_{0}\leftrightarrow{}^{3}P_{0}$ and the ${}^{1}S_{0}\leftrightarrow 4\text{f}^{13}6\text{s}^{2}5\text{d}\,(J=2)$ transitions in 171Yb. Thus, the measurement is susceptible to differential frequency shifts in the lasers used to drive these transitions. To suppress this noise source, the two lasers will be locked to the same optical cavity so that laser frequency drifts are common to the two transitions. The optical cavity can utilize crystalline reflective coatings Cole et al. (2016) to reduce the influence of thermal noise in the coatings on the effective cavity length.
@@
The accelerometric searches for scalar and vector dark matter [Fig. 3; sensitivity curves “JHU (DC)” and “MAGIS-100 (DC)” in Figs. 4 and 6, sensitivity curves “JHU (AC)” and “MAGIS-100 (AC)” in Fig. 6] are sensitive to differential forces between the two Yb isotopes. Based on previous atom-interferometric equivalence principle tests Asenbaum et al. (2020), the leading systematic effect of this kind is expected to arise from AC Stark shifts induced by the interferometry lasers. AC Stark shifts can be managed by utilizing a compensated optical spectrum Kovachy et al. (2015); Asenbaum et al. (2020) and by controlling the intensity, size, and divergence of the beams. We note that static equivalence principle tests are much more challenging from the perspective of systematic errors than searches for oscillating signals, as the static tests are subject to DC systematic shifts from magnetic fields, gravity gradients, and other sources Asenbaum et al. (2020).
@@
The largest source of systematic error for the axion search [Figs. 7 and 8; sensitivity curves “JHU (ST)” and “MAGIS-100 (ST)” in Fig. 9] will likely be the magnetic field, since the frequency of a $\Delta m=1$ transition is inherently sensitive to the magnetic field amplitude at first order. The magnetic field in the interferometer region will be controlled with a multiple-layer mu-metal shield Wodey et al. (2020), and the magnetic field along the interferometer trajectory will be measured in situ by using states with higher magnetic sensitivity Asenbaum et al. (2020).
@@
Several systematic effects will affect all three dark matter searches. For example, initial position and velocity displacements between interferometers produce phase shifts in the presence of gravity gradients. These phase shifts can be suppressed by overlapping the interferometer midpoints and by adjusting the frequencies of the interferometer lasers during the sequence Overstreet et al. (2018); Roura (2017). The Coriolis effect, which causes a velocity-dependent phase shift, can be reduced by counter-rotating the retroreflection mirror so that the interferometer laser direction remains constant in the freely falling frame of the atoms Dickerson et al. (2013); Lan et al. (2012). Shifts from blackbody radiation Haslinger et al. (2018) can be managed by controlling the temperature of the interferometry region.
@@
Ytterbium atom interferometry in atomic fountains can be used to perform sensitive searches for dark matter. The projected sensitivity to variations of the fine-structure constant is especially notable, as a differential frequency measurement between Yb electronic transitions is expected to surpass previous constraints on $d_{e}$ from Yb ion clocks by two orders of magnitude in the dark matter mass range from $10^{-22}$ eV to $10^{-16}$ eV. This improvement is made possible by combining the lower quantum projection noise of neutral-atom systems with the high sensitivity of open-f-shell electronic states to variations of the fine-structure constant. Compared to a proposed Yb optical clock Safronova et al. (2018), the JHU Yb atom interferometer is expected to have similar sensitivity while avoiding systematic effects from lattice light shifts and the density shift.
The most stringent search for vector dark matter with Yb atom interferometry can be derived from a static equivalence principle test between different atomic species, which is naturally sensitive to the static forces that would arise from a dark matter field sourced by the Earth. Since the vector coupling charge-to-mass ratios are similar for Yb isotopes and for the test masses used in the MICROSCOPE experiment Touboul et al. (2022), an improved search requires performing an atom-interferometric equivalence principle test with relative accuracy $\eta<10^{-15}$. Such a test is challenging but possible in laboratory-scale devices and is feasible with long-baseline experiments such as MAGIS-100. A more accurate equivalence principle test would also provide improved sensitivity to the scalar dark matter couplings parameterized by $d_{e}$, $d_{m_{e}}$, $d_{\hat{m}}$, and $d_{\delta m}$ Damour and Donoghue (2010). Searches for time-varying acceleration induced by vector dark matter can be implemented as well. In MAGIS-100, this type of measurement could enable an improved search for vector dark matter without requiring the control of DC systematic effects.
Atom interferometers can search for axion dark matter through the axion-fermion coupling, which induces spin torque. The 171Yb isotope, which has a ${}^{1}{S}_{0}$ electronic ground state and a nuclear angular momentum $I=1/2$ derived from a single unpaired neutron, is technically ideal for this experiment. We find that a spin torque search with 171Yb in MAGIS-100 would be competitive with existing limits from atomic magnetometers in some mass ranges.
@@ VI Acknowledgments
M.P. acknowledges support from a William H. Miller III Graduate Fellowship.
@@ VII Appendix
$Q_{e}=\left(-1.4+8.2\frac{Z}{A}+7.7\frac{Z(Z-1)}{A^{4/3}}\right)\times 10^{-4}.$ (15)
Similarly, when considering the scalar coupling to the electron mass, as mediated by the interaction term
@@
where $\lambda=h/(mc)$ is the Compton wavelength of the dark matter field and $d$ is the relevant coupling constant. In the limit $m\to 0$, the force becomes Coulomb-like. It can be shown Adelberger (2003) that the force on a test charge resulting from a uniformly-charged sphere of radius $R_{S}$ is
@@
$\begin{split}\tilde{q}_{\text{Earth}}&=q_{\text{Fe}}\Phi\left(\frac{R_{\text{Core}}}{\lambda}\right)\\ &+q_{\text{Si}}\left[\Phi\left(\frac{R_{\text{Earth}}}{\lambda}\right)-\Phi\left(\frac{R_{\text{Core}}}{\lambda}\right)\right]\\ &+q_{\text{O}}\left[\Phi\left(\frac{R_{\text{Earth}}}{\lambda}\right)-\Phi\left(\frac{R_{\text{Core}}}{\lambda}\right)\right].\end{split}$ (23)

### v3

@@ Abstract
Fig. 6 shows the projected sensitivity of atom-interferometric acceleration measurements to vector dark matter as a function of the dark matter particle mass. For the JHU projections, we assume the same experimental parameters as in Section III.A ($n=500$, maximum wave packet separation $2$ m). Likewise, the MAGIS-100 projections assume $n=500$ and $T=4.5$ s. Existing constraints from the MICROSCOPE Touboul et al. (2022), Eöt-Wash Shaw et al. (2022), and LIGO/VIRGO experiments Abbott (2022) are also shown. The searches for oscillating dark matter take into account the stochastic amplitude of the dark matter field Centers et al. (2021) by reducing the projected sensitivity by a factor of 3 throughout the mass range. These searches lose sensitivity at high frequencies when the period of the dark matter oscillation becomes shorter than the coherence time of a single measurement. Sensitivity at high frequencies can be enhanced by resonant detection schemes Graham et al. (2016b), which we do not consider here. In addition, the sensitivities for oscillating signals are based on a single-peak search algorithm. Using a multi-peak template Amaral et al. (2024) could provide higher sensitivity for some dark matter masses.
@@
The strongest searches for vector dark matter are expected to be derived from static equivalence principle tests. In this operating mode, the JHU apparatus is projected to reach a similar sensitivity to the MICROSCOPE experiment, while an analogous test in MAGIS-100 would be more sensitive by a factor of 10. Although the projected sensitivities of searches for time-oscillating forces are generally lower, we note that a search for oscillatory dark matter with ytterbium accelerometry in MAGIS-100 would reach a comparable sensitivity to MICROSCOPE and would be technically simpler than a static equivalence principle test. Compared to previously proposed searches for time-oscillating forces with strontium Yijun Jiang et al. (2021), we find that our approach would be a factor of 5 more sensitive, assuming identical interferometer parameters, due to the larger isotopic difference in $B-L$ and the use of a higher-frequency transition for the beamsplitter pulses.
III.3 C. Pseudoscalar dark matter
@@
Pseudoscalar particles (axion-like particles) are well-motivated dark matter candidates because they appear in theories that attempt to resolve other outstanding issues in fundamental physics, such as the strong CP problem Peccei and Quinn (1977) and the hierarchy problem Graham et al. (2015). Models inspired by string theory Arvanitaki et al. (2010) also predict the existence of low-mass axion-like particles. Here we consider the detection of such particles by means of the spin torque that they exert on nucleons. In the presence of a pseudoscalar field, a nuclear spin $\mathbf{s}$ experiences a Hamiltonian Graham et al. (2018)
@@
where $g_{\text{aNN}}$ is the coupling constant, the pseudoscalar field $a$ is given by
@@
the magnitude of the field $a_{0}=\sqrt{2\rho_{\text{DM}}}/(m_{\text{DM}}c)$, and the axion-like particle momentum is $\mathbf{p}$. This interaction, which has the same form as a magnetic interaction, gives rise to an energy shift between states with opposite spin direction. Assuming that the axion-like particle momentum is determined by its virial velocity, the energy shift is given by Graham et al. (2018)
@@
Figure 7: Interferometer geometry for searching for spin torque due to axion-like dark matter. Doppler-free two-photon transitions at time $t=0$ create superpositions of the ${}^{1}S_{0}$ electronic ground state and the ${}^{3}P_{0}$ excited state. The differential phase between the $\Delta m=1$ interferometer (a) and the $\Delta m=0$ interferometer (b) is sensitive to spin torque but insensitive to the electronic transition frequency. To reduce systematic effects, the two interferometers are spatially overlapped; interferometers are spaced vertically in the diagram for clarity.
@@
Figure 8: Alternative interferometer geometry for searching for spin torque due to axion-like dark matter. In (a), the atoms are initialized in the electronic ground state, and a Ramsey interferometer is implemented between hyperfine ground state levels. In (b), the Ramsey interferometer is instead carried out between magnetically sensitive states $\left|M\uparrow\right>$ and $\left|M\downarrow\right>$ (e.g. in the metastable ${}^{3}P_{2}$ level). The interferometer in (a) is sensitive to spin torque but relatively insensitive to the magnetic field, while the interferometer in (b) measures the magnetic field to characterize its systematic effect.
@@
Atom interferometers can be sensitive to this pseudoscalar dark matter coupling by searching for time dependence in the frequency of any transition that flips a nucleon spin. Following the proposal of Ref. Graham et al. (2018), an interferometer geometry that could be used for such a measurement is shown in Fig. 7. This configuration is sensitive to spin torque via the interferometer with the $\Delta m=1$ transition, while the interferometer with the $\Delta m=0$ transition is used to suppress systematic effects such as laser frequency drift. An alternative approach is illustrated in Fig. 8. Here the Ramsey interferometer between the hyperfine ground states of 171Yb is sensitive to the pseudoscalar field, while the second Ramsey interferometer between magnetically sensitive states (e.g. states in the metastable ${}^{3}P_{2}$ manifold) is used as a comagnetometer to decorrelate phase shifts from the magnetic field. We note that 171Yb, which has nuclear angular momentum $I=1/2$ from the spin of a single unpaired neutron, provides a simpler platform for this measurement than 87Sr Graham et al. (2018), which has $I=9/2$ from a combination of spin and orbital angular momentum.
@@
Figure 9: Constraint plot and projected sensitivities to axion-like dark matter coupled to nuclear spin. JHU (ST): projected sensitivity of spin torque search in JHU apparatus with interferometer geometry shown in Fig. 7 or Fig. 8. MAGIS-100 (ST): projected sensitivity of spin torque search with Yb in MAGIS-100. Also shown are existing constraints from atomic magnetometry experiments (Bloch et al. Bloch et al. (2020), green curve; Lee et al. Lee et al. (2023), purple curve; Wei et al. Wei et al. (2024), orange curve). Shaded regions are excluded by existing constraints. All curves are 95% confidence limits.
@@
The projected sensitivity of the JHU apparatus to axion-like dark matter through the axion-nucleon coupling is shown in Fig. 9. We also plot existing bounds from atomic magnetometer experiments Bloch et al. (2020); Lee et al. (2023) and the projected sensitivity of a spin torque search utilizing 171Yb in MAGIS-100, assuming the same phase resolution as in the JHU apparatus and an interferometer time of $T=4.5$ s. Although the laboratory-scale interferometer is not expected to be competitive with the magnetometer limits, a long-baseline interferometric measurement would have comparable sensitivity in several regions of parameter space. Our projections are more conservative than those reported in previous work Graham et al. (2018) due to our more conservative phase resolution estimates.
@@
The spectroscopic search for scalar dark matter [Figs. 1 and 2; projected sensitivity curve “JHU $\alpha(t)$” in Fig. 4] relies on a differential frequency measurement between the ${}^{1}S_{0}\leftrightarrow{}^{3}P_{0}$ and the ${}^{1}S_{0}\leftrightarrow 4\text{f}^{13}6\text{s}^{2}5\text{d}\,(J=2)$ transitions in 171Yb. Thus, the measurement is susceptible to differential frequency shifts in the lasers used to drive these transitions. To suppress this noise source, the two lasers will be locked to the same optical cavity so that laser frequency drifts are common to the two transitions. The optical cavity can utilize crystalline reflective coatings Cole et al. (2016) to reduce the influence of thermal noise in the coatings on the effective cavity length.
@@
The accelerometric searches for scalar and vector dark matter [Fig. 3; projected sensitivity curves “JHU (DC)” and “MAGIS-100 (DC)” in Figs. 4 and 6, projected sensitivity curves “JHU (AC)” and “MAGIS-100 (AC)” in Fig. 6] are sensitive to differential forces between the two Yb isotopes. Based on previous atom-interferometric equivalence principle tests Asenbaum et al. (2020), the leading systematic effect of this kind is expected to arise from AC Stark shifts induced by the interferometry lasers. AC Stark shifts can be managed by utilizing a compensated optical spectrum Kovachy et al. (2015); Asenbaum et al. (2020) and by controlling the intensity, size, and divergence of the beams. We note that static equivalence principle tests are much more challenging from the perspective of systematic errors than searches for oscillating signals, as the static tests are subject to DC systematic shifts from magnetic fields, gravity gradients, and other sources Asenbaum et al. (2020).
@@
The largest source of systematic error for the pseudoscalar search [Figs. 7 and 8; projected sensitivity curves “JHU (ST)” and “MAGIS-100 (ST)” in Fig. 9] will likely be the magnetic field, since the frequency of a $\Delta m=1$ transition is inherently sensitive to the magnetic field amplitude at first order. The magnetic field in the interferometer region will be controlled with a multiple-layer mu-metal shield Wodey et al. (2020), and the magnetic field along the interferometer trajectory will be measured in situ by using states with higher magnetic sensitivity Asenbaum et al. (2020).
@@
Several systematic effects will affect all three dark matter searches. For example, initial position and velocity displacements between interferometers produce phase shifts in the presence of gravity gradients. These phase shifts can be suppressed by overlapping the interferometer midpoints and by adjusting the frequencies of the interferometer lasers during the sequence. The idea of adjusting laser frequencies to compensate for vertical phase gradients, first proposed in Ref. Roura (2017), has been experimentally demonstrated in dual-isotope precision measurements Overstreet et al. (2018); Asenbaum et al. (2020). Crucially, this approach naturally improves as the sensitivity of the experiment improves, greatly relaxing constraints on how well the two isotopes must be overlapped spatially. Nevertheless, the initial position and velocity of each interferometer will need to be controlled at the level of $50~{}\mu$m and $50~{}\mu$m/s, respectively Asenbaum et al. (2020). The Coriolis effect, which causes a velocity-dependent phase shift, can be reduced by counter-rotating the retroreflection mirror so that the interferometer laser direction remains constant in the freely falling frame of the atoms Dickerson et al. (2013); Lan et al. (2012). We note that horizontal phase gradients from any source can be suppressed by appropriate changes to the angle of the retroreflection mirror during the interferometer sequence Asenbaum et al. (2020). As with vertical phase gradients, this compensation naturally improves with the sensitivity of the experiment. Shifts from blackbody radiation Haslinger et al. (2018) can be managed by controlling the temperature of the interferometry region.
@@
Ytterbium atom interferometry in atomic fountains can be used to perform sensitive searches for dark matter. The projected sensitivity to variations of the fine-structure constant is especially notable, as a differential frequency measurement between ytterbium electronic transitions is expected to surpass previous constraints on $d_{e}$ from ytterbium ion clocks by two orders of magnitude in the dark matter mass range from $10^{-22}$ eV to $10^{-16}$ eV. This improvement is made possible by combining the lower quantum projection noise of neutral-atom systems with the high sensitivity of open-f-shell electronic states to variations of the fine-structure constant. Compared to a proposed ytterbium optical clock Safronova et al. (2018), the JHU atom interferometer is expected to have similar sensitivity while avoiding systematic effects from lattice light shifts and the density shift.
The most stringent search for vector dark matter with ytterbium atom interferometry can be derived from a static equivalence principle test between different atomic species, which is naturally sensitive to the static forces that would arise from a dark matter field sourced by the Earth. Since the vector coupling charge-to-mass ratios are similar for ytterbium isotopes and for the test masses used in the MICROSCOPE experiment Touboul et al. (2022), an improved search requires performing an atom-interferometric equivalence principle test with a relative accuracy better than $10^{-15}$. Such a test is challenging but possible in laboratory-scale devices and is feasible with long-baseline experiments such as MAGIS-100. A more accurate equivalence principle test would also provide improved sensitivity to the scalar dark matter couplings parameterized by $d_{e}$, $d_{m_{e}}$, $d_{\hat{m}}$, and $d_{\delta m}$ Damour and Donoghue (2010). Searches for time-varying acceleration induced by vector dark matter can be implemented as well. In MAGIS-100, this type of measurement could enable an improved search for vector dark matter without requiring the control of DC systematic effects.
Atom interferometers can search for axion-like dark matter through the axion-fermion coupling, which induces spin torque. The 171Yb isotope, which has a ${}^{1}{S}_{0}$ electronic ground state and a nuclear angular momentum $I=1/2$ derived from a single unpaired neutron, is technically ideal for this experiment. We find that a spin torque search with 171Yb in MAGIS-100 would be competitive with existing limits from atomic magnetometers in some mass ranges.
@@ VI Acknowledgments
We thank Wei Ji, Jia Liu, and Ken Van Tilburg for useful discussions. M.P. acknowledges support from a William H. Miller III Graduate Fellowship.
@@ VII Appendix
$Q_{e}=\frac{Am_{N}}{m_{A}}\left(-1.4+8.2\frac{Z}{A}+7.7\frac{Z(Z-1)}{A^{4/3}}\right)\times 10^{-4}$ (16)
where $m_{N}$ is the mass of a nucleon and $m_{A}$ is the total mass of the atom. Similarly, when considering the scalar coupling to the electron mass, as mediated by the interaction term
@@
where $\lambda=h/(mc)$ is the Compton wavelength of the dark matter field and $d$ is the relevant coupling constant. In the limit $m\to 0$, the force becomes Coulomb-like. It can be shown Adelberger (2003); Hees (2018) that the force on a test charge resulting from a uniformly-charged sphere of radius $R_{S}$ is
@@
$\begin{split}\tilde{q}_{\text{Earth}}&=q_{\text{Fe}}\Phi\left(\frac{R_{\text{Core}}}{\lambda}\right)\\ &+q_{\text{Si}}\left[\Phi\left(\frac{R_{\text{Earth}}}{\lambda}\right)-\frac{V_{\text{Core}}}{V_{\text{Earth}}}\Phi\left(\frac{R_{\text{Core}}}{\lambda}\right)\right]\\ &+q_{\text{O}_{2}}\left[\Phi\left(\frac{R_{\text{Earth}}}{\lambda}\right)-\frac{V_{\text{Core}}}{V_{\text{Earth}}}\Phi\left(\frac{R_{\text{Core}}}{\lambda}\right)\right].\end{split}$ (24)